OpenStudy (anonymous):

Perform the division and explain your work in complete sentences: (2x+1)(3x+1)

OpenStudy (agent0smith):

What division? That's only multiplication.

OpenStudy (anonymous):

That's multiplication, but if you want the multiplication you use FOIL (first, outer, inner, last) so first you would multiply 2x by 3x which gets you 6x^2 because 2*3=6 and x*x=x^2. Now you do outer which is 2x*1 which equals 2x. Now inner is 3x*1 which is 3x and last is 1*1 which is 1. Now you have 6x^2+2x+3x+1 and now you simplify. You should get 6x^2+5x+1.